A Christmas tree ornament resembles a spherical convex mirror and is 8.7 cm in diameter. What is its focal length (in cm)?
Answer #1

focal length = R/ 2 = (8.7 / 2) / 2

f = 2.175 cm

=======

Comment before rate in case any doubt, will reply for sure.. goodluck